OpenStudy (anonymous):

how do you write (-8,0)(0,8) in slop intersept form

OpenStudy (anonymous):

y = mx + b, where m is your slope and b is your y intercept. to calculate m, simply find the ratio between the rise and run of your graph. \[m = (y2-y1)/(x2-x1) = (8-0)/(0-(-8)) = 8/8 = 1\] to find b, pick an ordered pair (x1,y1) use the formula \[y - y1 = m(x-x1) \rightarrow y - 0 = 1(x-(-8)) \rightarrow y = x + 8\] and the final answer is y = x - 8